OpenStudy (anonymous):

A recipe for chocolate chip cookies requires a half cup of sugar for a serving of 8 cookies. You need to make 30 cookies for your Math class and decide to use proportions to calculate how much sugar you need. Which of the following proportions is not set up correctly?

OpenStudy (anonymous):

8 cookies are used : For Half cup of Sugar, For 1 cookie, \(\frac{1}{8}\) half cups of sugar, Now for 30 cookies, \(\large \frac{1}{8} \times 30\) half cups of sugar.. Can you solve this??

OpenStudy (anonymous):

\[30 \; cookies = \frac{1}{8} \times 30 \times \frac{1}{2} = ??\]